solve by factoring check each answer
2x^2+3x=2
2x^2+3x-2=0
(2x-1)(x+2)=0
x=1/2
x=-2 this is how I did the problem but how do I SHOW TWO CHECKS

Answer by ankor@dixie-net.com(22740)

solve by factoring check each answer
2x^2+3x=2
2x^2+3x-2=0
(2x-1)(x+2)=0
x=1/2
x=-2
:
Substitute each solution in the original equation, see if it true
x=.5
2(.5^2) + 3(.5) = 2
2(.25) + 1.5 = 2
.5 + 1.5 = 2
and
x=-2
2(-2^2) + 3(-2) = 2
2(4) - 6 = 2
8 - 6 = 2